### v2

# Language Models are Crossword Solvers

Abstract
Crosswords are a form of word puzzle that require a solver to demonstrate a high degree of proficiency in natural language understanding, wordplay, reasoning, and world knowledge, along with adherence to character and length constraints. In this paper we tackle the challenge of solving crosswords with Large Language Models (LLMs). We demonstrate that the current generation of state-of-the art (SoTA) language models show significant competence at deciphering cryptic crossword clues, and outperform previously reported SoTA results by a factor of 2-3 in relevant benchmarks. We also develop a search algorithm that builds off this performance to tackle the problem of solving full crossword grids with LLMs for the very first time, achieving an accuracy of 93% on New York Times crossword puzzles. Contrary to previous work in this area which concluded that LLMs lag human expert performance significantly, our research suggests this gap is a lot narrower.
Language Models are Crossword Solvers
Soumadeep Saha  and Sutanoya Chakraborty  and Saptarshi Saha  and Utpal Garain Indian Statistical Institute Kolkata, India Correspondence: soumadeep.saha97@gmail.com
1 Introduction
Crossword puzzles are a type of word game that typically takes the form of a square grid of white and black squares. The objective of the puzzle is to fill the white squares with letters from words or phrases based on the provided clues (see Figure 1). Although crosswords come in a variety of styles, the two most popular ones are the American style or straight crosswords and cryptic crosswords.
Figure 1: Example of a crossword puzzle (left) and cryptic clues (right). (left) The grid must be filled up with answers from the semantic clues provided. The gray highlighted squares produce additional constraints, e.g. first character of answer to clue 1 (across) must be the same as the first character of answer to clue 4 (down). Example by Fred Piscop. (right) In cryptic crosswords, the clues involve some form of wordplay and synonyms, and often involve world knowledge. Examples are taken from the cryptonite dataset.
Modern LLMs demonstrate astounding skills in reasoning, coding, wordplay, question answering (QA), and a multitude of other tasks (Wei et al., 2022a; Minaee et al., 2024). In this paper we attempt to analyze LLMs’ ability to solve crossword puzzles, with the primary goal of understanding strengths and weaknesses demonstrated by SoTA LLMs.
Solving crossword puzzles requires high levels of proficiency in understanding contextual clues, wordplay, world knowledge, and reasoning (see Figure 1 (right)). To illustrate the kinds of skills involved, consider as an example the cryptic crossword clue - ‘‘Culminating point of story about Judy’s husband by the railway track (5,4)’’.  [^1] Example taken from https://lovattspuzzles.com. Solving this clue requires connecting ‘Judy’ to the popular puppet show Punch and Judy, and inferring that ‘Judy’s husband’ refers to PUNCH. Additionally, we must observe that ‘railway track’ is synonymous to LINE, and combining these gives PUNCHLINE which also means ‘Culminating point of story’.
There is another aspect to solving crossword puzzles, which is constraint satisfaction (see Figure 1). Not only does the generated output have to adhere to the length limitations, it also has to incorporate character level constraints from previously generated answers. Successful crossword solvers generate answers based on clues, length information, character information from earlier answers, etc. and have the ability to backtrack when faced with inconsistent solutions in a targeted manner. Thus, even when the crossword clues are simpler, as in straight crosswords, the problem still poses a significant challenge to SoTA LLMs. Since solving crosswords requires proficiency in several desirable areas, our objective with this paper is to probe LLMs’ abilities at this multi-faceted task.
To that end, we first perform extensive analysis to understand how well LLMs can solve crossword clues based on the provided contextual information and constraints. We show that current-generation SoTA LLMs demonstrate massively improved performance when compared to previous SoTA baselines, often outperforming them by a factor of 2 or 3, without any fine-tuning.
In our analysis, we found that the ability of LLMs to adhere to length constraints, which is a critical component of solving crossword puzzles, is somewhat limited. We present results demonstrating that an LLM’s ability to count the number of constituent characters in a word degrades with the prevalence of the word. This suggests that their ability to perform this task does not generalize well.
Armed with the knowledge gathered from our analysis, we tackle the full problem - i.e. not only answering clues, but filling in the grid considering all the imposed constraints. We device an algorithm called SweepClip which can solve straight crossword puzzles with an accuracy of 93.1% with the aid of an LLM. This algorithm exploits the constraints imposed by previously generated answers to improve future answers, and achieves results comparable to rule-based systems. This algorithm also improves the clue deciphering capability of LLMs, more than doubling the initial performance.
To check if the performance improvements shown by LLMs is due to dataset contamination, we perform tests with post knowledge-cutoff benchmarks and human evaluation. We believe our work has been a major step towards solving crosswords with LLMs, and with future advances in certain key areas, LLMs’ performance in this task will be comparable to, if not better than human experts.
2 Background
Traditional approaches to solving crossword puzzles typically involve two parts - a candidate answer proposal system and a grid-filling algorithm (Angelini et al., 2005; Ginsberg, 2011). One such system, Proverb, Littman et al. (2002) achieves a letter accuracy of $98.1\%$ on crosswords from the New York Times (NYT). The candidate answer proposal systems rely on large datasets of crossword clues and answers (Shazeer et al., 1999) and grid-filling relies on algorithms for the constraint satisfaction problem (Angelini et al., 2005; Zugarini et al., 2023; Shazeer et al., 1999; Kulshreshtha et al., 2022). Kulshreshtha et al. (2022) who curated a dataset of NYT crossword puzzles and established benchmarks with language models have highlighted this task as “… a new high bar for AI systems”.
Attempts to solve cryptic crossword clues with traditional algorithms, employing a large dataset of clues and a curated context free grammar (CFG) based parser (Deits, 2015), were not found to be performant ($7\%$ accuracy according to Sadallah et al. (2024)). Using LLMs to solve cryptic crossword clues has received attention in some recent papers. Efrat et al. (2021) curated a large dataset of cryptic crossword clues sourced from two UK dailies - The Times and The Telegraph and fine-tuned a T5-Large (Raffel et al., 2020) model to establish baseline performances. They proposed a training split for their dataset where the answers in the training and test sets are mutually exclusive so the model cannot rely on memorization. Following this, we have the work by Rozner et al. (2021), where they curated a dataset of cryptic crossword clues from The Guardian, and fine-tuned a T5-Large model with a curriculum learning approach, demonstrating performance improvements. Rozner et al. (2021) had some salient criticisms of the work by Efrat et al. (2021), and suggested that a disjoint train-test split is not a strong enough condition to ensure that the models learn to solve cryptic crosswords, since the models demonstrate “…robustness to plural and other inflections”. Rozner et al. (2021) proposed a split where words with similar roots are grouped together, and demonstrated that this stringent criteria leads to reduced performance.
The most recent work by Sadallah et al. (2024) presents results with recent LLMs like Mistral-7B (Jiang et al., 2023), LLaMA2-7B (Touvron et al., 2023) and ChatGPT  [^2] It is unclear to us which version is used.(OpenAI, 2021) in few-shot settings and also by fine-tuning the Mistral model. They report that ChatGPT surpasses other models with an accuracy of $9.5\%$, which demonstrates a significant performance gap with respect to human experts, who solve $99\%$ of cryptic crossword clues (Sadallah et al., 2024). They point out key limitations in their work, like the extremely limited set of LLMs used, and potential for data contamination.
All three recent works on cryptic crosswords conclude that there is a massive gap in performance between LLMs and human experts. Also, these recent works focusing on solving cryptic crosswords with LLMs do not attempt to solve the full problem, i.e. they do not attempt to incoporate grid information and the constraints imposed by them, and rather elect to pose the problem of solving crosswords in the form of a question answering (QA) task. Kulshreshtha et al. (2022) attempted to solve NYT crossword puzzles with language models and a Satisfiability Modulo Theories (SMT) solver with limited success, and had to resort to culling the crossword grid based on candidate generations and ground truth answers.
In this paper, we undertake the challenge of solving a full crossword puzzle by developing an algorithm that incorporates grid information based on the answers given by an LLM. We find that with the application of our algorithm, the base performance of LLMs can be boosted significantly, and that the gap between the performance of LLMs and human experts is much smaller than previously anticipated.
3 Analyzing Crossword Clue Solving
In this section we analyze the performance of several LLMs at different scales at the clue deciphering task.
3.1 Datasets
Most of our analysis is performed on three crossword puzzle datasets. The first one covering straight or American style crosswords was curated by us from the very popular and long running puzzle section of the NYT. The other two, namely, Cryptonite by Efrat et al. (2021) and word-init-disjoint by Rozner et al. (2021) cover cryptic crossword puzzles. As discussed in Section 2, the two cryptic crossword datasets have methodological differences in how they partition training-test splits, however, this is not a very pertinent consideration for us, since we do not perform any training. A bulk of the results are reported on the NYT dataset we curated for straight crosswords and word-init-disjoint (abbreviated as Init) for cryptic crosswords since it was found to be more challenging than Cryptonite for LLMs. See Appendix B for more details.
3.2 Answering Crossword Clues
The first step in solving a crossword puzzle is deciphering its clues, so following in the scheme of Rozner et al. (2021); Sadallah et al. (2024); Kulshreshtha et al. (2022) and Efrat et al. (2021) we first explore this as a QA task, where the clue and its length is provided to the LLM as a query, with the expectation that it will produce the corresponding answer.
We report results with Phi 3 3.8B Instruct (Abdin et al., 2024), Mistral 7B Instruct (Jiang et al., 2023), LlaMa 2 70B (Touvron et al., 2023), LlaMa 3 8B Instruct, Llama 3 70B (Meta, 2024), Mixtral 8x7B (Jiang et al., 2024), Claude 3 Sonnet (Anthropic, 2024), GPT 3.5 Turbo, and GPT 4 Turbo (OpenAI, 2023) to cover a wide range of parameter scales and a mix of open-weights and proprietary models. We investigate the performance of these models with few-shot prompts (5-shot, 10-shot) on samples from the NYT dataset (ours), Cryptonite and the word-init-disjoint split (see Figure 2).
Figure 2: Analyzing LLMs’ ability to generate answers from crossword clues. We test LLMs at different scales on the NYT, Cryptonite and word-init-disjoint dataset with 5-shot, 10-shot prompts. We show that current generation of LLMs show significantly improved performance, and in particular GPT 4 Turbo outperforms the previous SoTA by 2$\sim$ 2$\times$ 2 × on the word-init-disjoint dataset with 5-shot prompts.
We find that the performance difference between 5-shot and 10-shot  [^3] We elected not to perform experiments with 10-shot prompts on Claude and GPT-4-Turbo due to budget constraints. prompted answers is not significant (see Figure 2), which is consistent with the findings by Sadallah et al. (2024). We also note that there is an appreciable difference (up to 5%) in performance of models across the board between Cryptonite and Init, with LLMs showing diminished performance on the Init dataset.
The models demonstrate improved performance with scale across datasets, and in particular show impressive performance on the NYT dataset with Llama 3 70B, GPT 3.5 Turbo, Claude 3 Sonnet, and GPT-4-Turbo achieving 27.2%, 26.05%, 37.7% and 41.2% accuracy (exact match) respectively.
The performance of LLMs on cryptic crosswords is worse compared to straight crosswords, however, Claude 3 Sonnet and GPT-4-Turbo outperform previous SoTA results on both the cryptic crossword datasets, achieving an accuracy of 12.9%, 23.5% on Cryptonite and 10.8%, 18.7% on Init respectively. The performance of GPT-4-Turbo is rather extraordinary, with a 1.97$\times$ improvement over previous SoTA (9.5% reported by Sadallah et al. (2024), see Figure 2). LlaMa 3 70B and GPT 3.5 Turbo demonstrate performance comparable to rule-based systems, and the smaller LLMs we tested perform far worse.
Further details about experimental protocols can be found in Appendix C and the details of the results summarized in Figure 2 can be found in Appendix F.
3.3 Exploiting Partially Filled Grids
In the course of solving a crossword puzzle, we encounter intermediate states where some of the clues have been deciphered. Successful crossword solvers exploit these clues in order to inform their decisions about future answers. For example, in Figure 1 (left), when we want to solve for the clue in position 14 (across), we can use the characters from position 2 (down) and 4 (down) to narrow down the set of possible answers to only those that fit the template “_ T _ P”. In this section we study LLMs’ performance at exploiting these constraints to arrive at better answers (see Table 1).
For this experiment we chose the best performing open-weights and proprietary model i.e. LLaMA 3 70B and GPT-4-Turbo respectively. We also report results on some smaller models to see if the trends hold. We report results on the NYT dataset and Init dataset with 5-shot prompts. For each query, $k$of the characters (letters) of the answer is provided alongside the clue and expected length of the answer, and the LLMs are expected to “unmask” the remaining characters using the provided constraint information and the crossword clue.  [^4] Further experimental details can be found in Appendix C.
Hint (%) 0% 25% 50% 70%
Model NYT init NYT init NYT init init
Mistral 7B (5-shot) 10.95% 1.70% 9.70% 2.80% 11.95% 4.80%
LlaMa 3 8B (5-shot) 15.8% 1.30% 19.7% 2.85% 24.65% 6.25%
LlaMa 3 70B (5-shot) 27.20% 6.40% 31.80% 11.45% 45.30% 20.35%
GPT 4 Turbo (5-shot) 41.2% 18.70% 59.95% 33.70% 75.75% 52.85% 76.30%
Mistral 7B (fine-tuned) 27.0%
Table 1: LLMs can exploit information provided by solving other clues to improve their accuracy. During crossword solving, we encounter partially filled grids, which provides hints about potential solutions in the form of character constraints. We test LLMs to see if they can incorporate this information and demonstrate improved performance. $\dagger$ is the results reported by Sadallah et al. (2024) by fine-tuning a Mistral 7B model on cryptic crossword solving. We demonstrate that models can indeed exploit these constraints and we find that GPT-4-Turbo with 5-shot prompts outperforms the fine-tuned Mistral 7B model by a factor of $\sim$2.8$\times$.
We observe (Table 1) that, in all but one case, LLMs show improved performance with an increasing percentage of constraint information for both datasets. Additionally, to compare the performance of GPT-4-Turbo to previously reported SoTA results by Sadallah et al. (2024) we perform the experiment with the same settings and dataset split as them, to find that GPT-4-Turbo with 5 shot prompts (76.3% accuracy) outperforms the fine-tuned Mistral 7B model (27% accuracy) by a factor of 2.8$\times$.
The fact that LLMs can successfully exploit constraints to answer crossword clues better, suggests that they are well suited to the task of solving full crosswords. The massive jump in performance observed in modern SoTA LLMs at this task and the task in Section 3.2 is extremely serendipitous, and we investigate this in further detail in Section 5.3.
3.4 Sub-token Counting
Despite massive performance gains demonstrated by current SoTA LLMs, one area in which they perform poorly is adherence to length constraints, suggesting they lack the ability to perform sub-token counting i.e. counting the number of characters present in a word or phrase. We observed that even the best performing model (GPT-4-Turbo) produces answers of incorrect length on 26.2% and 16.9% of the Init and NYT datasets respectively.
The fact that LLMs are stumped by this rather simple task is on the surface a little baffling, however, we must consider that most SoTA LLMs first tokenize sequences using some Byte-Pair Encoding (Sennrich et al., 2016) like scheme. Thus, during the word embedding stage in Transformers (Vaswani et al., 2017), when the token is converted to an embedding vector, information about constituent characters in the token is lost, and must be regained somehow during training. Although we are not sure how LLMs regain this length information during training, we have a hunch.
There are websites  [^5] https://word.tips/words-by-length/ for example. which contain large lists of words with their corresponding lengths. Often replies in message boards also include a count of the number of characters in the reply. Artifacts like these, which contain enough information to infer length of tokens, go on to become part of the datasets that LLMs are trained on. We hypothesise that LLMs learn to count sub-tokens based on this information provided during training.
To investigate this further, we devised the sub-token counting task, wherein the LLM is provided a sequence of (lower case) characters without white-spaces and asked to predict the number of characters making up the sequence. To test our hypothesis, we consider three sets of 1000 (English) words - Common, Medium and Rare based on word unigram frequencies curated by Segaran and Hammerbacher (2009) from Google’s Trillion Words corpus. The Common, Medium and Rare words have ranks in the range of 1-5,000, 47,500-52,500 and 95,000-100,000 respectively.
Figure 3: LLMs ability to count the number of characters in a word decline with frequency of the word.
If our hypothesis is incorrect, and language models have a widely generalizable ability to perform sub-token counting, we should see no difference in counting performance across words with different prevalence. However, we observe that (see Figure 3) the accuracy of LLMs at the sub-token counting task declines with frequency of the token, for all LLMs tested. We further analyzed if there was a difference in sub-token counting performance between words that are part of the model vocabulary and randomly generated gibberish following the same distribution of lengths, to account for potential distribution shifts between frequent and rare words. To do this we first created a set of words by taking an intersection of all words that are part of every open source model in consideration and the list of top 100,000 words. This is to ensure that the sequences are - (a) extremely likely [^6] We can’t be sure for the proprietary models to be vocabulary tokens for every model in consideration and (b) are not special tokens like <bos>. Then we created a set of gibberish words by replacing each character of the vocab set words with a randomly chosen character from the set {a-z}, thus guaranteeing that they have the same length distributions.
Model Vocab. Gibberish
[…]
Phi 3 3.8B Instruct $79.4$ $61.2$
Mistral 7B Instruct $47.9$ $28.2$
Llama 3 8B Instruct $92.6$ $69.7$
Mixtral 8x7B $92.6$ $80.1$
Llama 2 70B $92.8$ $80.0$
Llama 3 70B $99.6$ $87.5$
GPT 3.5 Turbo $86.0$ $62.1$
GPT 4 Turbo $99.8$ $98.8$
Table 2: LLM sub-token counting performance for vocabulary words and gibberish. In addition to declining counting performance with decreasing prevalence of the word, there is a sizeable difference in counting performance between vocabulary words and randomly generated gibberish.
We find (see Table 2) that in addition to counting accuracy being affected by frequency, there is often a huge disparity between the accuracy of vocabulary vs gibberish words. Although, this does not conclusively show that LLMs rely on memorized training instances to perform sub-token counting, it does provides strong evidence suggesting that LLMs learn to count based on length information containing artifacts in training data. Future works exploring this idea further would be compelling.
4 SweepClip - Our proposed algorithm to solve crosswords with LLMs
Emboldened by our successes in deciphering answers from crossword clues with LLMs, we set out to address the challenge of tackling the problem of filling crossword grids with LLM assistance. Note that this task not only involves generating correct answers from provided clues, but hinges on exploiting constraints from already generated words, and back-tracking to eliminate past generations that do not fit well when new evidence becomes available. Since LLMs demonstrate this ability, when paired with the right search algorithm, it should be possible to solve to crosswords with the aid of LLMs.
Algorithm 1 - SweepClip (Sketch)
1:Given crossword grid, clues and an LLM.
2:Generate candidate answers for all clues. (sweep)
[…]
4:     Prune answers not in largest connected component. (clip)
5:     Prune conflicting answers.
6:     Generate answers for neighbors of surviving answers.
7:     if solved or budget_exceeded then
8:         break
9:     end if
10:end for
Our proposed algorithm (outlined in Algorithm 1 and detailed in Appendix A) first generates a set of candidate answers for all clues provided with the crossword (sweep) and uses a graph based criteria to eliminate answers that do not fit well to the grid (clip). Following this, we use the constraints generated from the previous step answers, generate more candidate answers, and prune the bad fitting candidates. We iteratively apply this strategy, until either - (i) the entire crossword is filled, (ii) we exceed a preset number of iterations, or (iii) we run out of LLM computational budget.
For pruning we use the largest connected component from the answers generated so far to ensure that the generated answers are somewhat coherent amongst themselves. This is over-restrictive, as it is possible that isolated answers are correct, however, we find that this strict pruning strategy works better to eliminate bad answers early rather than using the constraints imposed by potentially bad answers to generate further bad answers.
We would like to highlight that our algorithm is sub-optimal, and does not provide convergence or correctness guarantees beyond elimination of conflicting answers. Note that this problem is NP-Hard and it is difficult to find approximate solutions for it (Kulshreshtha et al., 2022). This algorithm was designed from the ground up to minimize cost incurred from calls to the LLM by avoiding branching based on LLM answers.
5 Results
In this section we present results from our algorithm at the crossword solving task and present further results investigating the performance demonstrated by current generation SoTA LLMs.
5.1 Solving Cryptic Crossword Clues
We observed a massive performance increase in cryptic crossword clue deciphering in the current generation of SoTA LLMs. We present comparisons between the results we obtained on the cryptic crossword clue answering task 777treated as QA with the Init dataset, and previously reported SoTA results in Table 3.
Model Method Acc.
Rule-based CFG+WordNet
(Deits, 2015) 7.3
T5 (Efrat et al., 2021) SFT 1.1
T5 Curriculum
(Rozner et al., 2021) Learning 6.5
Mistral 7B SFT 1.2
Mistral 7B 10 shot 4.6
Chat GPT 3 shot 9.5
(Sadallah et al., 2024)
GPT 4 Turbo (ours) 5 shot 18.70
GPT 4 Turbo (ours) CoT(1)@3SC 20.85
Table 3: Comparison of our results with previously reported SoTA Results.
We also note that Chain-of-thought (Wei et al., 2022b) prompting (1 shot) with self-consistency Wang et al. (2023) (3 samples) leads to further performance gains. Our best result on the word-init-disjoint split is 20.85% which improves over the previous SoTA (9.5%) by a factor of 2.2$\times$ without any fine-tuning.
5.2 Solving Crossword Puzzles
For this task we employed our algorithm SweepClip on a set of 100 randomly sampled Monday NYT crossword puzzles. We used two LLMs for this - GPT-4-Turbo with max_depth of 30 and a budget of 0.5 USD per crossword, and LLaMA 3 70B with max_depth of 35 and a budget of 600 LLM calls.
Error Tolerance % of Crosswords
LLaMa 3 GPT-4 T
100% solved 0 48
$\leq$ 1 character error 1 55
$\leq$ 5 character error 10 71
$\geq 90\%$ Accuracy 30 80
$\geq 50\%$ Accuracy 82 98
Table 4: Results from solving NYT crosswords with our algorithm SweepClip.
We find (see Table 4) that our algorithm with GPT-4-Turbo solves 48% of crosswords without any errors, and 55% of crosswords with at most 1 character wrong. The average character level accuracy in crossword solving is 93.1% ($\pm$ 14.1%). Our algorithm improves the clue-wise answer accuracy 888clue level accuracy is different from character level accuracy (Kulshreshtha et al., 2022), e.g. it is possible to have a filled in crossword without deciphering all clues. (exact match) to 89.6% ($\pm$ 16.9%) from the base accuracy (without the algorithm) which is 43.5% ($\pm$ 23.5%) an improvement of 2.1$\times$. Note the previously reported SoTA accuracy on this task was 26% with retrieval augmented generation and an SMT solver coupled with an oracle that eliminates parts of the crossword grid that do not have suitable generated answers (Kulshreshtha et al., 2022).
The performance for the smaller LLaMA 3 70B is worse, however, our algorithm still manages to improve final clue answering accuracy to 59.4% ($\pm$ 24.1%) from a base accuracy of only 22.3% ($\pm$ 14.4%).
Our algorithm implicitly performs self-consistency checks to improve candidate answers. For example the generated clue at the first sweep may be discarded and regenerated with a different prompt and then accepted if it does not produce conflicts, and is coherent with a larger number of other answers. Thus with the application of our algorithm, we have successfully exploited the constraint information to boost the performance of LLMs far beyond what would have been possible with straightforward QA like clue deciphering. To the best of our knowledge this is the first algorithm that demonstrates succesfully solving crosswords with the aid of LLMs.
5.3 Data Contamination and Generalizability
Having seen remarkable performance improvements in the current generation of SoTA LLMs, we set out to investigate the reason behind this success. To rule out data-contamination, we curated additional cryptic crossword clue datasets comprising entirely of puzzles published after May 20, 2024, which is after the knowledge cutoff date of all LLMs examined. These datasets are sourced from The Guardian [^9] https://www.theguardian.com/ and Lovatts Puzzles [^10] https://lovattspuzzles.com. Note, the Init dataset is also sourced from the The Guardian, thus the results reported in Table 5 should be consistent.  [^11] More details on data curation is in Appendix B.
Model Lovatts Guardian init
Acc. Acc. Acc.
LlaMa 3 70B 26.03% 5.5 % 6.4 %
Claude 3 Sonnet 46.28% 12.5% 10.8%
GPT 4 Turbo 61.57% 18.5% 18.7%
Table 5: Performance of LLMs on curated datasets that appeared after knowledge cut-off. Note, Init by Rozner et al. (2021), also sourced their data from The Guardian, thus these results provide a fair head-to-head comparison of performance.
We see no appreciable difference in performance on the post-cutoff dataset, (see Table 5) leading us to conclude that these LLMs can generalize beyond potential contamination in their training set. The performance of models on the Lovatts Puzzles dataset (post-cutoff) is even better than that on the Guardian dataset.
To ascertain if the models can reason about unseen cryptic crossword clues to arrive at correct responses, we elect to perform human evaluation121212 Further details in Appendix E of model responses. We employ a 3-shot Chain-of-thought prompt to elicit a reasoned response to crossword clues from GPT-4-Turbo, which are then analyzed by our team for soundness vis-à-vis factual and logical errors. We choose 100 samples from the post-cutoff Lovatts set and rely on consensus of all evaluators to ascertain if a response contains factual or logical errors (e.g. wrong counting, wrong anagram, drawing a conclusion that does not follow from the premise).
Sound Unsound
(61%) (39%)
Correct (65%) 48% 17%
Wrong (35%) 13% 22%
Table 6: Results from human evaluation of reasoning provided by GPT-4-Turbo. An answer is called correct if the final answer provided by the model matches the ground truth exactly. The answer is called sound if it contains no logical or factual errors.
The results (see Table 6) show that 74% of the time GPT-4-Turbo provided a correct answer, it also gave sound reasoning in support of the answer. This leads us to conjecture that they possibly have a significant ability to reason and generalize.
6 Conclusions and Discussions
Solving crosswords requires proficiency in a multitude of desirable skills, and cryptic crosswords were previously thought to be firmly in the domain of human experts. Our findings challenge this notion. We have demonstrated that the current generation of SoTA LLMs show significantly improved aptitude at solving straight and cryptic crossword puzzle clues, and can exploit constraints provided by partially solved crosswords to boost this performance further. We also found that this emergent ability generalizes well to the post knowledge cut-off regime and is accompanied with the capacity to produce reasoned explanations.
We have also developed an algorithm that, with the aid of LLMs, further boosts this performance, and achieves 93% accuracy on Monday New York Times crosswords. With the aid of this algorithm it is possible to achieve massive performance gains even if the baseline LLM accuracy is low, thus indicating that, when paired with the right search strategy, LLMs can succesfully solve crosswords.
This calls into question previously reported results which claim that LLMs massively lag behind human experts at solving cryptic crosswords. We are firmly of the opinion that the current gap between human experts and LLMs will cease to exist in the near future. Incorporating length constraints effectively still remains a challenge, and further enhancements in numerical reasoning abilities in LLMs is a promising avenue to pursue.
Limitations
Limitation 1 - Reporting crossword solving results We only reported results on solving Monday New York Times crossword puzzles and didn’t report Tuesday-Thursday which typically increase in level of difficulty. We also did not solve cryptic crossword grids. There are primarily two reasons for this. (i) As we have seen in the paper, even if the base LLM has relatively low accuracy when paired with a search/prune algorithm, the accuracy can be improved considerably. There is a trade-off between base accuracy and the number of LLM calls, i.e. if the base accuracy is lower, the algorithm needs to make a lot more calls to the LLM increasing computational costs. We have observed that it is possible to solve harder NYT (straight) crosswords or cryptic crosswords, but at an increased computational budget. This makes reporting statistics cost-prohibitive for us at this time. (ii) The cryptic crossword datasets do not come with grid information, and instead provide cryptic clues and answers as question-answer pairs. To successfully demonstrate cryptic crossword grid solving we would need to acquire a high quality cryptic crossword dataset with full grid information.
Limitation 2 - Algorithm Our algorithm is potentially sub-optimal as it discards potential correct answers and does not fully explore the consequences of each possible generated answer. This is again a computational cost saving technique. Ideally, we would generate one candidate answer, followed by all its neighbors, and whenever there is a conflict, branch out and explore all options to see which is a better fit. We elect not to do this, because it involves a potentially exponential number of LLM calls. Note that this problem is NP-Hard, however, future studies with much larger computational budget for LLM calls will definitely have more success employing a more thorough search strategy.
Limitation 3 - Sub-token Counting We have demonstrated that LLMs show limitations at sub-token counting, and have hypothesised that they do not have widely generalizable skill at this task, rather relying on memorization. The evidence provided in this paper is compelling, but definitely not enough to conclusively prove this hypothesis. Further studies are required that can intervene at the pre-training stage of LLMs to conclusively demonstrate whether this hypothesis is True. We do not possess the computational power to undertake such a study at this stage.
Limitation 4 - Fine-tuning We do not report results with fine-tuned models. It is possible that further improvements could be seen with fine-tuning, however, we are interested in studying the emergent properties of large language models in this context instead of building special purpose models that can solve crosswords.
Ethics Statement
In keeping with ACL ethical guidelines we make all scientific artifacts generated for this study freely available and open source under the MIT licence, this includes all software we created, every prompt that was used, all raw model outputs, and data we generated. We do not distribute the New York Times crossword dataset since it is the intellectual property of The New York Times. We have purchased a subscription to access their puzzle repository, but used the data without their consent for this study. The use of their data falls within the terms of the Fair Use doctrine set within 17 U.S.C. §107. To aid in reproducibility we provide a list of dates to uniquely identify the crossword puzzles used in the study, and provide detailed instructions on how to acquire them through the New York Times.
We foresee no serious ethical implications on society at large from this study.
Appendix A Appendix - Details of the Sweep Clip Algorithm
This section is intended to provide details of the SweepClip algorithm introduced in Section 4.
A crossword puzzle consists of a grid (see Figure 1) and a set of clues $C=\{c_{1},\ldots c_{n}\}$ and answers (ground truth) $A=\{a_{1},\ldots a_{n}\}$ corresponding to position $i$ in the grid. The grid imposes a graph $G=(V,E)$, where $V=\{v_{1},\ldots v_{n}\}$ is the set of vertices corresponding to every clue/answer in the crossword, and
$\begin{split}E=\{(v_{i},v_{j})|&\;\forall i,j\;i\neq j\>s.t.\;a_{i},a_{j}\textrm{ share a}\\ &\textrm{common grid position}\}\end{split}$ (1)
Given an LLM, and a set of clues $C^{\prime}\subseteq C,C^{\prime}=\{c_{j_{1}},c_{j_{2}},\ldots\}$ corresponding to vertices $\{v_{j_{1}},v_{j_{2}},\ldots\}$ we can generate candidate answers $\hat{A}=\{\hat{a}_{j_{1}},\hat{a}_{j_{2}},\ldots\}$, where $\hat{a}=LLM(c)$. We abbreviate this as $\hat{A}=LLM(C^{\prime})$.
Two candidate answers $\hat{a}_{i},\hat{a}_{j}$ are said to be in conflict, if $\mu$-th position of $a_{i}$ and $\nu$-th position of $a_{j}$ are in the same grid position, however $\hat{a}_{i}[\mu]\neq\hat{a}_{j}[\nu]$, i.e. the $\mu$-th character of $a_{i}$ and $\nu$-th character of $a_{j}$ are different.
There are two sub-graphs $G_{p}(\hat{A})$ and $G_{n}(\hat{A})$ of $G$ that correspond to a set of candidate answers $\hat{A}$. Let $v_{i},v_{j}$ correspond to $\hat{a}_{i},\hat{a}_{j}\in\hat{A}$. An edge $(v_{i},v_{j})\in E$ is in $G_{p}(\hat{A})$ if and only if $\hat{a}_{i},\hat{a}_{j}$ don’t conflict, else it is in $G_{n}(\hat{A})$. We call the largest connected component of a graph $H$, $LCC(H)$, and the neighborhood of a set of vertices $V$ in the graph $G$, $ngbd(G,V)$. The algorithm is detailed in Algorithm 2.
Algorithm 2 - SweepClip (Detailed)
1:Given $C$, crossword graph $G$ and an LLM.
2:Generate $\hat{A}=LLM(C)$.
[…]
4:     Construct $G_{p}(\hat{A}),G_{n}(\hat{A})$
5:     $L:=LCC(G_{p}(\hat{A}))$.
6:     $\hat{A}=\{\hat{a}_{i}|\;v_{i}\in L\}$
7:     while $G_{n}(\hat{A})$ has edges. do
8:         Max degree vertex in $G_{n}(\hat{A})=v_{m}$
9:         Remove $v_{m}$ from $G_{n}(\hat{A})$.
10:         $\hat{A}=\hat{A}-\hat{a}_{m}$
11:     end while
12:     Calculate $N=ngbd(G,\hat{A})$.
13:     $\hat{A}=\hat{A}+LLM(N)$
14:     if solved or budget_exceeded then
15:         break
16:     end if
17:end for
Appendix B Appendix - Datasets
Dataset (Clues) Train Validation Test
Cryptonite 470,804 26,156 26,157
word-init-disjoint 75,847 32,628 33,905
NYT (Clues) 10,000 2,000
NYT (Grids) 100 (7700 clues) Test only
After May 20, 2024
Lovatts 242
The Guardian 200
Table 7: Dataset details
This section contains details of the datasets used, and details relating to how they were curated.
B.1 Cryptonite and word-init-disjoint
Cryptonite is a cryptic crossword clue dataset introduced by Efrat et al. (2021), and is gathered from The Times and The Telegraph - two popular UK dailies. In their official split, the training and test sets do not contain any common answers, so that a model trained on the train split cannot rely on memorization to answer the cryptic crossword clue. Two thousand random samples samples were chosen from the train split to report results in Section 3. The Cryptonite dataset [^13] https://github.com/aviaefrat/cryptonite is not distributed under an open source licence.
word-init-disjoint is also a cryptic crossword clue and was introduced by Rozner et al. (2021). This is gathered from The Guardian, and their official split introduces the additional constraint that if two answers share a common root (first two characters) they are placed in the same split. The results reported in 3 are based on two thousand randomly sampled clues from their official test split. We distribute all materials pertaining to this dataset [^14] https://github.com/jsrozner/decrypt with this paper.
B.2 New York Times Dataset
We collected and curated two datasets from the New York Times  [^15] https://www.nytimes.com/crosswords/ for our analysis of straight crossword puzzles. The first consists of 100 randomly sampled Monday crossword puzzles ranging from 20th January 1969 to 7th August 2023 with all clues (7700) and grid information, which were used to report results in Section 5.
Additionally, we curated 12,000 randomly sampled clue answer pairs split into two sets - test (2000) and support (10,000). All three sets are completely disjoint. [^16] https://www.xwordinfo.com/ is useful for older NYT crosswords
The test and support sets are distributed alongside this paper in the Supplementary Materials, however, the 100 complete grids are not, as they are intellectual property of the New York Times. They can be accessed after purchasing a subscription to their service. To aid reproducibility we include the list of dates to uniquely identify the crossword grids used to report results.
B.3 Post-cutoff datasets
We curated two post-cutoff datasets to test for potential dataset contamination. One of them was sourced from Lovatts puzzles  [^17] https://lovattspuzzles.com/ and The Daily Mail. In total 244 clues and answers were manually curated from puzzles published strictly after May 20, 2024. To maintain parity with word-init-disjoint dataset by Rozner et al. (2021) we also obtained 200 clue answer pairs from The Guardian puzzles published after May 20, 2024.
We cross-checked both these datasets for duplicate clues against all 665,497 clues from the Cryptonite and word-init-disjoint datasets. We found two duplicates for the Lovatts set which were removed. The dataset obtained from The Guardian contained no duplicates. This does not guarantee that these cryptic crossword clues have not appeared in any form previously, but we estimate that probability to be low, since there were very few duplicates when matched against the extremely large cryptic crossword datasets. Both these datasets are distributed alongside the paper in the Supplementary Materials.
B.4 Sub-token counting
For the task in Section 3.4 we first obtained a list of 300,000+ words and their unigram frequencies provided by Segaran and Hammerbacher (2009) and sourced from the Google Trillion Words corpus. For the Rare, Medium and Common words we randomly selected 1000 words each from ranks 95,000-100,000, 47,500-52500 and 0-5000 respectively. The Vocab set was constructed by taking an intersection of the vocabulary tokens of every open source language model in consideration, i.e. Phi 3, LLaMA 3 70B, LLaMA 3 8B, LLaMA 2, Mistral 7B, and Mixtral. Further, we took the intersection of this set with the set of top 100,000 words to make sure the chosen tokens are actual words and not special tokens. We sampled 1000 words from the resulting set to create the vocab set. For the gibberish set, we replaced each character from each word in every set with a character randomly sampled from the set {a-z} and randomly sampled 1000 words from it. This ensures that the distribution of lengths of words seen in the gibberish set is close to the distribution of lengths in the other sets. These 5 sets are distributed alongside the paper in the Supplementary Materials.
Appendix C Appendix - Experimental Details
C.1 General Overview
Model Params. Context Knowledge
Length Cut-off
Phi 3 mini Instruct 3.8B 4K Oct. 2023
Mistral v0.2 Instruct 7B 32K Dec. 2023
LLaMA 3 Instruct 8B 8K Mar. 2023
Mixtral v0.1 8x7B 32K Dec. 2023
LLaMA 2 70B 4K Sep. 2022
LLaMA 3 70B 8K Dec. 2023
Claude 3 sonnet 20140229 ? 200K Mar. 2024
GPT-3.5-Turbo-0125 ? 16K Sep. 2021
GPT-4-Turbo 2024-04-09 ? 128K May 2024
Table 8: Models used and their details.
We used temperature=0.5 for all models and experiments. For few-shot prompt responses max-tokens were set to 10, and for Chain-of-thought responses max-tokens was set to 1000. top_p, top_k or any other parameters were not changed. The Jinja templates used to format the text prompts are provided alongside the paper, and are borrowed from the model-cards of the respective models. Claude and GPT models were accessed through their respective APIs and all other models were run locally on one server consisting of 2x80GB Nvidia A100 GPUs. All models were used in bf16 format whenever supported. Format of prompts are given in Appendix D.
C.2 Experimental Protocols for Section 3.3
[…]
$\begin{split}&h=\\ &\max\Bigg{(}1,\text{round}\Big{(}\frac{k}{100}\times len(\text{answer})\Big{)}\Bigg{)}\\ &\forall k>0\end{split}$ (2)
$h$ many characters are randomly selected and revealed, all other characters are replaced with “_”.
Appendix D Appendix - Prompts
In this section we will detail the prompts used for all experiments presented in the paper. Every prompt used in generating outputs presented in this paper is provided alongside the paper in Supplementary Materials.
D.1 Few Shot Prompts
For the experiments presented in Section 3.2 we used prompts in the following format -
    "role":"system", "content":
    "You are an expert crossword
    solver. Given a clue please
    provide the best possible
    answer succinctly. Do not
[…]
    number of characters in the
    answer is given in brackets
    and must be strictly adhered
    to. e.g. Clue: Daily update
    (4)// means the answer should
    have 4 characters."
   "role":"user", "content":
[…]
    Clue: <query clue> (length) // "
For the experiments presented in Section 3.3 we used prompts in the following format (with the same system prompt) -
"role":"user", "content":
"Clue: <clue 1> (3) // _ _ N
[…]
Clue: < ...  > (6) //
[…]
Clue: <query clue> (length) //
            _ _ X _ X => "
D.2 Chain-of-thought Prompts
For the results presented in Table 3 and for human-evaluation we used Chain-of-thought prompts. These were based on 10 examples we manually created and are diverse in terms of skills required to solve them. An example Chain-of-thought reasoning prompt is as follows -
"Clue: ecstasy in old forbidden city,
[…]
Typically a cryptic crossword
has synonyms and wordplay
[…]
’forbidden city’ is a part
[…]
can also mean the drug MDMA
[…]
’ecstasy in old forbidden city’
[…]
e in Peking can be peeking
which is a synonym of
[…]
peeking is 7 characters and
therefore fits the character
[…]
Appendix E Appendix - Human Evaluation and Reasoning Examples
Human evaluation was performed by 3 of the authors of this paper. They were asked if a particular response contained logical or factual errors and provided detailed instructions and examples of the kinds of errors to be wary of. All responses were evaluated by at least 2 authors and if there were conflicting answers, discussions were held to reach a consensus. All evaluated responses alongside human-annotated labels are provided alongside this paper.
In this section we provide a few Chain-of-thought elicited reasoning examples given by GPT-4-Turbo on the Lovatts dataset. These are sampled from the set used to perform human evaluation.
1. Clue: Initially, any region extending around land
Answer (Ground Truth): AREA
GPT-4-Turbo Output:
’Initially’ is often used in cryptic crosswords to indicate taking the initial letters of the following words.
’any region extending around land’ gives us the words Any, Region, Extending, and Land.
Taking the initial letters of each word, we get A, R, E, and A.
Putting these letters together forms the word AREA.
’Area’ is a synonym for ’region’ and fits the clue both in definition and letter count.
AREA is 4 characters and therefore fits the character count.
Answer: AREA
2. Clue: Fur stayed in audition for medical help
Answer (Ground Truth): FIRSTAID
GPT-4-Turbo Output:
’Fur’ can mean pelt, coat, hide.
’stayed’ can mean remained, lingered.
’in audition’ indicates a homophone clue (sounds like).
’for medical help’ can mean aid, care, assistance.
Combining these:
- ’Fur stayed in audition’ suggests a homophone of a phrase where ’fur’ and ’stayed’ combine to sound like another word.
- ’Fur’ as ’fur’ and ’stayed’ sounds like ’aid’ (homophone of stayed).
- Putting them together doesn’t form a clear word, so reconsidering ’fur’ might lead to another homophone, such as ’first aid’ (fur sounded + stayed sounding like first aid).
’First aid’ fits as a term for ’medical help’.
’First aid’ is 8 characters and therefore fits the character count.
Answer: FIRST AID
3. Clue: Be earlier because of self-esteem
Answer (Ground Truth): FOREGO
GPT-4-Turbo Output:
’Be earlier’ can suggest synonyms such as precede, antedate.
’because of’ is often used in cryptic clues as a linking phrase or to indicate causation.
’self-esteem’ can be linked to synonyms like pride, ego.
Breaking down the phrase:
’Be earlier’ suggests a word meaning to come before - "precede" fits but is too long.
Let’s consider "ego" for ’self-esteem’, which is a common synonym.
Looking at ’Be earlier because of self-esteem’, we can interpret this as a word that indicates being earlier due to ego, or ego causing one to be first or ahead.
The answer "PRE" (as in prefix, or before) combined with "EGO" gives us PREGO, which is 5 letters and does not fit. However, considering synonyms and length, the correct answer could be "PRECEDE" if we think of it as "PRE" (before) + "CEDE" (yield, coming from ego or self-esteem yielding to others). But this is 7 letters.
Reevaluating, if we simplify to just consider "ego" leading or being earlier, we can think of "PREGO" as a playful or informal way to say "before" due to ego, but the character count is off.
The best fit for a 6-letter word considering the clue and synonyms would be "BEFORE," which directly means "be earlier" and can be influenced by one’s self-esteem (ego) driving their actions to be ahead or first. This interpretation fits the clue structure and the character count.
Answer: BEFORE
Appendix F Appendix - Results
Model Common Medium Rare
Phi 3 66.9% 58.8% 55.7%
Mistral 7B 38.9% 34.6% 27.2%
Llama 3 8B 81.7% 78.1% 75.2%
Mixtral 8x7B 79.1% 74.5% 74.0%
Llama 2 70B 85.3% 78.6% 76.1%
Llama 3 70B 93.5% 91.3% 90.7%
GPT 3.5 Turbo 84.2% 78.8% 76.4%
GPT 4 Turbo 99.3% 94.3% 92.7%
Table 9: Data for results plotted in Figure 3
Model Method Accuracy (NYT) Accuracy (Crypt) Accuracy (Init)
Phi 3 3.8B Instruct 5 shot 5.15 0.80 0.35
10 shot 5.55 0.90 0.30
Mistral 7B Instruct 5 shot 11.50 1.65 1.70
10 shot 11.60 2.05 2.00
Llama 3 8B Instruct 5 shot 15.90 1.80 1.30
10 shot 14.45 2.20 1.30
Mixtral 8x7B 5 shot 12.70 3.85 2.65
10 shot 13.30 4.65 2.10
Llama2 70B 5 shot 17.75 2.95 2.85
10 shot 18.40 3.95 2.35
Llama3 70B 5 shot 27.20 7.85 6.40
10 shot 28.20 8.50 5.85
GPT 3.5 Turbo 5 shot 26.05 7.90 5.55
10 shot 26.65 7.95 6.40
Claude 3 Sonnet 5 shot 37.70 12.9 10.8
GPT 4 Turbo 5 shot 41.20 23.5 18.7
Table 10: Data for results plotted in Figure 2